# 1]AgiBot2]Shanghai AI Lab3]CUHK 4]SJTU 5]FDU 6]HKUST 7]HIT ]

### v2

@@ Abstract
We introduce EnerVerse, a generative robotics foundation model that constructs and interprets embodied spaces. EnerVerse employs an autoregressive video diffusion framework to predict future embodied spaces from instructions, enhanced by a sparse context memory for long-term reasoning. To model the 3D robotics world, we propose Free Anchor Views (FAVs), a multi-view video representation offering flexible, task-adaptive perspectives to address challenges like motion ambiguity and environmental constraints. Additionally, we present EnerVerse-D, a data engine pipeline combining the generative model with 4D Gaussian Splatting, forming a self-reinforcing data loop to reduce the sim-to-real gap. Leveraging these innovations, EnerVerse translates 4D world representations into physical actions via a policy head (EnerVerse-A), enabling robots to execute task instructions. EnerVerse-A achieves state-of-the-art performance in both simulation and real-world settings
@@ 1 Introduction
Creative AI in vision has achieved significant progress, especially in video generation, where models produce high-quality videos from human instructions Kong et al. (2024); Zheng et al. (2024). This success highlights the model’s spatiotemporal imagination, enabling accurate forecasting of future frames. Similarly, robotic manipulation, a fundamental task in embodied AI, needs accurate prediction of future actions based on language instructions to interact with the physical world. Based on this sharing principle of future space prediction, an natural strategy is to align robotics action prediction with a video generation task to leverage video generation models’ imagination capabilities for policy planning. Motivated by this, recent studies Wen et al. (2024); Rigter et al. (2024); Cheang et al. (2024); Guo et al. (2024) have conducted preliminary explorations by fine-tuning general video generation models on robotic manipulation videos to align feature representations with the robotics domain, and predict physical actions. However, such methods Rigter et al. (2024) often simply adapt general-purpose video generation models to embodied tasks, neglecting the substantial gap between their representation space and the three-dimensional, temporally interconnected robotics environment, thereby hindering accurate action policy prediction.
@@
To bridge the gap, we propose EnerVerse, a generative robotics foundation model designed to construct and interpret the robotics world. In EnerVerse, we employ an autoregressive video diffusion framework that iteratively predicts the embodied future space based on a given instruction. Within this generative paradigm, we define a minimal unit of the future space as a ‘chunk’, and the model repeatedly predicts the next chunk to incrementally expand the space. Additionally, to prevent model collapse and enhance the action planning capabilities, we design a sparse context memory mechanism during training. Instead of relying on consecutive memory, this mechanism preserves essential prior content throughout the generation process in a non-redundant manner, theoretically allowing infinite-length sequence generation. This autoregressive strategy and the sparse memory design enable stable 2D embodied video generation, yet bridging the approach to 3D robotics remains challenging.
To address this, we introduce the novel concept of Free Anchor View (FAV) in EnerVerse, by simultaneously predicting multi-view robotics videos to represent the 3D robotics world. Unlike hardware-mounted cameras that provide fixed perspectives, FAVs are not physically affixed but freely positioned by the users, offering flexible, task-adaptive multi-view perspectives as shown in Fig. 1. This flexibility is particularly beneficial for robotic tasks, as it mitigates challenges such as: (1) motion modeling ambiguity caused by extrinsic changes in body-mounted cameras during whole-body motions; (2) physical constraints in confined spaces, like collisions in narrow kitchens; and (3) the limited generalization and adaptability of static workspace-mounted cameras. Multi-FAV setups provide richer visual information, constructing implicit 3D spatial representations that significantly improve policy performance.
@@
Figure 1: An overview of EnerVerse. Leveraging camera observations, we first obtain a 3D reconstruction via depth warping and rendering Lassner and Zollhofer (2021), then setup FAVs. EnerVerse that (a) connects to a video generator head (EnerVerse-G) to produce multi-view videos, (b) attaches to a robotic action policy head (EnerVerse-A) for action prediction, and (c) integrates with 4DGS to form a data flywheel (EnerVerse-D) for Sim2Real adaptation.
@@
However, acquiring precisely calibrated multi-camera observations, along with robotic actions, is costly and labor-intensive, limiting the EnerVerse’s access to large-scale real-world training data. Simulators can generate abundant synthetic data, but the sim-to-real gap remains a significant challenge. To overcome this, we propose EnerVerse-D, a data engine that combines a generative model with 4D Gaussian Splatting (4DGS). By leveraging the generative model’s adaptability and 4DGS’s spatial constraints, EnerVerse-D initiates a self-reinforcing data flywheel, narrowing the sim-to-real gap and enhancing EnerVerse ’s performance.
@@
The contributions of this work are as follows: (1) We develop an innovative chunk-wise autoregressive diffusion model architecture capable of logically reasoning for embodied future space, benefiting from a sparse contextual memory mechanism. (2) We propose a novel FAV-based embodied future space method associated with policy planning, significantly enhancing spatial understanding. (3) We construct a data flywheel in the robotics domain that integrates 4DGS optimization into multi-view video generation, enabling iterative improvements in data quality.
2 Related work
Video Generation Models. Diffusion-based video generation models have made notable progress, especially in text-to-video (T2V) generation Blattmann et al. (2023); Song et al. (2020). Early T2V approaches Zhang et al. (2023); Chen et al. (2023); Ren et al. (2024); Guo et al. (2023) build on text-to-image (T2I) priors by introducing temporal modules trained on video data. DynamicCrafter Xing et al. (2025) reuses motion priors from T2V diffusion models in an image-to-video (I2V) context. Recent works  Kong et al. (2024); Zheng et al. (2024); Bao et al. (2024) explores replacing U-Nets with Diffusion Transformer (DiT) Peebles and Xie (2023a). Other studies Gao et al. (2024) incorporate causal mechanisms to generate longer sequences or extend video-generation models into world modeling by forecasting future states Hu et al. (2023); Bruce et al. (2024); Wang et al. (2023). In this paper, we adopt DynamicCrafter as our base I2V framework due to its open-source availability and widespread use. We also ensure compatibility with modern DiT architectures, although that is not our main focus here.
Video Pretraining for Robotics. GR-2 Cheang et al. (2024) presents a generalizable robot manipulation framework that pretrains on large-scale internet videos, then fine-tunes on both video generation and action prediction for robotic trajectories. LAPA Ye et al. (2024) uses non-robot action videos for representation learning, mapping discrete latent actions (via VQ-VAE) to robotic manipulation tasks through a VLA model. SEER Tian et al. (2024) further explores inverse dynamics pretraining to boost performance. AVID Rigter et al. (2024) employs DynamicCrafter Xing et al. (2025) as its foundation, using an adapter for the robotics domain. VidMan Wen et al. (2024), based on OpenSora Zheng et al. (2024), focuses on environment prediction before action generation but is limited to 2D image space. In contrast, we propose generating long-sequence futures via a novel data generation engine, capturing richer motion information vital for robotics.
4D Generation. Recent progress Chen and Wang (2024) allows reconstruction of dynamic scenes from 2D videos using 3D GS (GS) Kerbl et al. (2023) and Neural Radiance Fields (NeRF) Mildenhall et al. (2021). Prior approaches approximate the spatio-temporal 4D volume with sets of 4D Gaussians Yang et al. (2023), jointly optimizing geometry and motion in canonical space Wu et al. (2024a). More recent advancements Li et al. (2024) employ customized sampling for multi-view video diffusion models, particularly for single dynamic objects. DimensionX Sun et al. (2024) leverages multiple LoRAs Hu et al. (2021) for diverse camera motions, while Cat4D Wu et al. (2024b) uses a single multi-view diffusion model to generate videos for dynamic 3D reconstruction. By contrast, our method produces videos from a Free Anchor View tailored for robotic manipulation tasks. In our offline data flywheel stage, GS complements video-generation models to mitigate the Sim2Real gap.
3 Methods
Figure 2: An overview of our chunk-wise autoregressive generation approach. (a) During training, random clean frames selected from consecutive sequences are combined with noisy frames to predict denoised latents. (b) In the inference phase, once new denoised frames are generated, they become the next set of clean frames for subsequent inference steps. This iterative procedure continues until the EoS frame is detected. For clarity, we illustrate only a single view of the autoregressive process, although multi-view generation is fully supported by the model.
EnerVerse comprises several designs, including a chunk-wise autoregressive generation framework and the FAV design for embodied future space generation. We additionally integrate a 4DGS to construct a data flywheel, referred to as EnerVerse-D, and a policy head to generate physical actions, referred as EnerVerse-A.
@@ 3.1 Next Chunk Diffusion
Chunk-wise Autoregressive Generation. As shown in Fig. 2, the observed latent sequence is represented as $\bm{o}_{t}^{1:K}=[\bm{o}_{t}^{1},\dots,\bm{o}_{t}^{K}]\in\mathbb{R}^{K\times H\times W\times C}$, encoded by a pre-trained Variational Autoencoder (VAE). Here, $K$ denotes the number of observed frames, $H\times W$ represents the spatial resolution, $C$ is the number of channels, and $t$ is the denoising step. Similarly, the latent representation of the rendered image is given by $\bm{r}_{t}^{1:J}\in\mathbb{R}^{J\times H\times W\times C}$. For simplicity, we treat $\bm{r}$ as a special case of $\bm{o}$. The predicted latent sequence is denoted as $\bm{z}_{t}^{1:M}=[\bm{z}_{t}^{1},\dots,\bm{z}_{t}^{M}]\in\mathbb{R}^{M\times H\times W\times C}$. The goal is to develop a video diffusion model that generates these predicted latents conditioned on $\bm{o}_{0}^{1:K}$ and a textual prompt $\bm{c}$, following the conditional probability: $p_{\theta}(\bm{z}_{t}^{1:M}\mid\bm{c},\bm{o}_{t}^{1:K})$. Here, $\theta$ represents the parameters of the denoising network, which is defined as $\bm{\epsilon}_{\theta}(\bm{z}_{t}^{1:M},\bm{c},\bm{o}_{t}^{1:K},t)$. The network is trained to predict the ground truth noise $\mathbf{\epsilon}$ from the noisy frame targets by optimizing the loss function:
@@
where $\mathbf{\epsilon}$ is the sampled ground truth noise, and $\theta$ denotes the learnable parameters. Consistent with prior work Salimans and Ho (2022), we predict $v$ in practice. After training, the denoised data $\mathbf{z}_{0}$ can be derived from random noise $\mathbf{z}_{T}$ through iterative denoising.
@@
Context Frame Mechanism. Instead of the conventional approach of using consecutive frames as the clean frame context for chunk prediction during training, we propose using sparsely sampled frames as the clean frame context. This approach leverages the redundancy often present in video data, allowing approximately 80% of frames to be discarded without compromising training effectiveness. Additionally, the high frame-dropping ratio enhances the model’s robustness, particularly in handling out-of-distribution (OOD) scenarios such as covariant shift problems commonly encountered in the robot learning domain. From a representation learning perspective, this randomized sampling strategy promotes a deeper understanding of chunk prediction, potentially outperforming methods that rely on continuous frames.
@@
Figure 3: Multi-View Diffusion Generator Block. Observational frames from Camera $i$ or Anchor View $i+1$ is encoded using a VAE and the corresponding ray direction maps are then concatenated with the video latents. Subsequently, a combination of convolutional layers and attention mechanisms is applied. Notably, the observational frames from the camera are optional.
Single-view or fixed multi-anchor view Goyal et al. (2024) approaches face challenges in handling occlusion and physical constraints in complex 3D environments, such as kitchens, where fixed cameras may be obstructed or infeasible. Similarly, wrist-mounted cameras complicate policy learning by coupling environmental dynamics with the robot’s motion. To address these limitations, we extend the diffusion generator described in Sec. 3.1 to a Free Anchor View video generation pipeline.
As illustrated in Fig. 3, our method directly generates multi-view latents, represented as $\bm{z}_{t}^{1:M}\in\mathbb{R}^{M\times V\times H\times W\times C}$, where $V$ denotes the number of views. To ensure consistency across different views, a ray direction map, encoding intrinsic and extrinsic camera parameters, is concatenated with the corresponding image latents along the channel dimension, enabling view-aware generation through ray casting chen2023ray; sajjadi2022scene. Additionally, spatial attention is applied along the $V$ dimension to model cross-view relationships, ensuring coherent and consistent outputs while preserving the geometric relationships among objects in the scene.
@@
Figure 4: The pipeline for EnerVerse as a data engine. Observation images from multiple cameras and rendered anchor view images are processed by the multi-view video generator to produce denoised multi-view videos. These videos, along with their camera poses, are used in 4DGS for 4D scene reconstruction. The reconstructed 3D content is rendered into anchor views to generate high-precision images. These high-quality rendered images are iteratively refined and fed back into the pipeline, improving motion consistency and reconstruction accuracy. This iterative process ensures geometric consistency and produces high-fidelity outputs suitable for applications.
Real-World Data Flywheels. Collecting calibrated multi-camera observations in real-world settings is expensive and labor-intensive. Consequently, we rely on data from simulators, which often exhibit significant domain gaps when applied to real-world scenarios. These gaps, including discrepancies in visual appearance and metric accuracy, hinder their direct applicability. To address these issues, we propose a multi-stage data generation pipeline that utilizes sparse observations to generate multi-view videos of a scene as shown in Fig. 4. First, a base model EnerVerse is fine-tuned to be capable of receiving a complete offline observation sequence. When inferring, observations captured from multiple mounted cameras, covering robotic arm motion and scene dynamics to ensure cross-view consistency, are used to construct a 4D Gaussian representation via GS. Once the 3D scene reconstruction is complete, the anchor views are rendered to obtain higher-precision, geometrically consistent observations. These rendered observations are iteratively refined with EnerVerse-D to generate pseudo-ground truth. After collecting sufficient real-world multi-view video data, we further fine-tune the multi-view video generator. This iterative process reduces noise, improves reconstruction accuracy, and facilitates Sim2Real domain adaptation, ultimately producing large-scale, high-quality datasets critical for training 4D generation models.
Figure 5: Visualization of FAVs on the LIBERO. Anchor View 1 represents the observation image captured by a mounted camera. Anchor View 2 and Anchor View 3 are generated by rendering from a point cloud reconstructed from Anchor View 1 using depth wrapping.

### v3

# EnerVerse: Envisioning Embodied Future Space for Robotics Manipulation

@@ Abstract
We introduce EnerVerse, a generative robotics foundation model that constructs and interprets embodied spaces. EnerVerse employs a chunk-wise autoregressive video diffusion framework to predict future embodied spaces from instructions, enhanced by a sparse context memory for long-term reasoning. To model the 3D robotics world, we adopt a multi-view video representation, providing rich perspectives to address challenges like motion ambiguity and 3D grounding. Additionally, EnerVerse-D, a data engine pipeline combining generative modeling with 4D Gaussian Splatting, forms a self-reinforcing data loop to reduce the sim-to-real gap. Leveraging these innovations, EnerVerse translates 4D world representations into physical actions via a policy head (EnerVerse-A), achieving state-of-the-art performance in both simulation and real-world tasks. For efficiency, EnerVerse-A reuses features from the first denoising step and predicts action chunks, achieving about 280 ms per 8-step action chunk on a single RTX 4090. Further video demos, dataset samples could be found in our project page.
@@ 1 Introduction
Creative AI in vision has achieved significant progress, especially in video generation, where models produce high-quality videos from human instructions kong2024hunyuanvideo ; opensora . This success highlights the model’s spatiotemporal imagination, enabling accurate forecasting of future frames. Similarly, robotic manipulation, a fundamental task in embodied AI, needs accurate predictions of future actions based on language instructions to interact with the physical world. Based on this sharing principle of future space prediction, one natural strategy is to align robotics action prediction with a video generation task to leverage video generation models’ imagination capabilities for policy planning. Motivated by this, recent studies wen2024vidman ; rigter2024avid ; cheang2024gr ; guo2024prediction have conducted preliminary explorations by fine-tuning general video generation models on robotic manipulation videos to align feature representations with the robotics domain, and predict physical actions. However, such methods rigter2024avid often simply adapt general-purpose video generation models to embodied tasks, neglecting the substantial gap between their representation space and the three-dimensional, temporally interconnected robotics environment, thereby hindering accurate action policy prediction. We do not claim a direct monotonic link between pixel-level video quality and control success. Rather, we align the latent space to encode 3D, action-conditioned dynamics so that actions can reliably follow generated trajectories.
@@
To bridge the gap, we propose EnerVerse, a generative robotics foundation model designed to construct and interpret the robotics 4D (3D with time) world. In EnerVerse, we employ an autoregressive video diffusion framework that iteratively predicts the embodied future space based on a given instruction. Within this generative paradigm, we define a minimal unit of the future space as a ‘chunk’, and the model repeatedly predicts the next chunk to incrementally expand the space. Additionally, to prevent model collapse and enhance the action planning capabilities, we design a sparse context memory mechanism during training. Instead of relying on consecutive memory, this mechanism preserves essential prior content throughout the generation process in a non-redundant manner, theoretically allowing infinite-length sequence generation. While this design achieves stable 2D embodied video generation, it remains insufficient for 3D understanding.
@@
Figure 1: An overview of EnerVerse. With camera observations, we first obtain a 3D reconstruction via depth warping, then multi rendered images. EnerVerse (a) connects to a video generator head (EnerVerse-G) to produce multi-view videos, (b) attaches to a robotic action policy head (EnerVerse-A) for action prediction, and (c) integrates with 4DGS to form a data flywheel (EnerVerse-D) for Sim2Real.
A straightforward solution is directly extending 2D video generation into multi-view video generation and mounting multiple cameras to provide more 3D cues goyal2024rvt . However, adding cameras increases hardware costs, I/O bandwidth requirements, and system complexity. To circumvent these challenges, we argue that a strong 3D generative prior learned during pre-training can effectively enhance the single-camera setup. During inference, observations from the single camera can be wrapped and rendered into multiple views. To establish this generative prior, we pre-train EnerVerse on multi-view video generation and introduce a Multi-View Diffusion Generator Block. This block utilizes ray direction maps to encode camera information and employs temporal attention to seamlessly fuse multi-camera data. Emperically, pretraining with multi-view consistency supplies a 3D prior that benefits even single-camera deployments via rendered auxiliary views.
@@
Although the pre-training stage requires substantial multi-camera data, acquiring precisely calibrated multi-camera observations with corresponding robotic actions is costly and labor-intensive. While simulators can generate abundant synthetic data, the Sim2Real gap remains a significant challenge. To address this, we propose EnerVerse-D, a data engine combining a generative model with 4D Gaussian Splatting (4DGS). By leveraging the adaptability of the generative model and the spatial constraints of 4DGS, EnerVerse-D establishes a data flywheel that narrows the Sim2Real gap.
@@
The contributions of this work are as follows: (1) A chunk-wise autoregressive diffusion architecture with sparse contextual memory capable of long-term grounding. (2) A multi-view diffusion generator that endows the model with a 3-D spatial prior beneficial under single-camera deployment. (3) A 4DGS-based data flywheel that supplies geometry-consistent multiview training data for robotics.
2 EnerVerse
Figure 2: An overview of our chunk-wise autoregressive generation approach and multi-view diffusion generator block. (a) During training, random clean frames from consecutive sequences are combined with noisy frames to predict denoised latents. In inference, newly generated denoised frames become the next clean frames for subsequent steps, iterating until the EoS frame is detected. Only a single view of the autoregressive process is shown for clarity. (b) In the multi-view diffusion generator block, observational frames from Camera $i$ or Rendered View $i+1$ are encoded with a VAE. Ray direction maps are concatenated with video latents, followed by conv layers and attention mechanisms.
EnerVerse comprises several designs, including a chunk-wise autoregressive generation framework and the sparse memory design for embodied future space generation. We additionally integrate a 4DGS to construct a data flywheel, referred to as EnerVerse-D, and a policy head to generate physical actions, referred as EnerVerse-A.
@@ 2.1 Next Chunk Diffusion
Chunk-wise Autoregressive Generation. As shown in Fig. 2, the observed latent sequence is represented as $\bm{o}_{t}^{1:K}=[\bm{o}_{t}^{1},\dots,\bm{o}_{t}^{K}]\in\mathbb{R}^{K\times H\times W\times C}$, encoded by a pre-trained Variational Autoencoder (VAE). Here, $K$ denotes the number of observed frames, $H\times W$ represents the spatial resolution, $C$ is the number of channels, and $t$ is the denoising step. Similarly, the latent representation of the rendered image is given by $\bm{r}_{t}^{1:J}\in\mathbb{R}^{J\times H\times W\times C}$. For simplicity, we treat $\bm{r}$ as a special case of $\bm{o}$. The predicted latent sequence is denoted as $\bm{z}_{t}^{1:M}=[\bm{z}_{t}^{1},\dots,\bm{z}_{t}^{M}]\in\mathbb{R}^{M\times H\times W\times C}$. The goal is to develop a video diffusion model that generates these predicted latents conditioned on $\bm{o}_{0}^{1:K}$ and a textual prompt $\bm{c}$, following the conditional probability: $p_{\theta}(\bm{z}_{t}^{1:M}\mid\bm{c},\bm{o}_{t}^{1:K})$. Here, $\theta$ represents the parameters of the denoising network, which is defined as $\bm{\epsilon}_{\theta}(\bm{z}_{t}^{1:M},\bm{c},\bm{o}_{t}^{1:K},t)$. $\bm{c}$ is encoded by a frozen T5 encoder and then projected with an MLP. The network is trained to predict the ground truth noise $\mathbf{\epsilon}$ from the noisy frame targets by optimizing the loss function:
@@
where $\mathbf{\epsilon}$ is the sampled ground truth noise, and $\theta$ denotes the learnable parameters. We follow salimans2022progressive to implement the v-prediction. Instead of predicting the noise $\epsilon_{t}$, the model predicts $v_{t}$, defined as: $v_{t}=\alpha_{t}\epsilon_{t}-\sigma_{t}x_{0}$. Here, $\alpha_{t}=\sqrt{\bar{\alpha}_{t}}$ (signal scale) and $\sigma_{t}=\sqrt{1-\alpha_{t}^{2}}$ (noise scale), consistent with the forward process equation $x_{t}=\alpha_{t}x_{0}+\sigma_{t}\epsilon_{t}$. After training, the denoised data $\mathbf{z}_{0}$ can be derived from random noise $\mathbf{z}_{T}$ through iterative denoising.
@@
Sparse Memory Mechanism. Instead of the conventional approach of using consecutive frames as the clean frame context for chunk prediction during training, we propose using sparsely sampled frames as the clean frame context. This approach leverages the redundancy often present in video data, allowing approximately 80% of frames to be discarded without compromising training effectiveness. Additionally, the high frame-dropping ratio enhances the model’s robustness, particularly in handling out-of-distribution (OOD) scenarios such as covariant shift problems commonly encountered in the robot learning domain. From a representation learning perspective, this randomized sampling strategy promotes a deeper understanding of chunk prediction, potentially outperforming methods that rely on continuous frames.
@@
Single-view video generation struggles to recover accurate 3D structure and resolve occlusions, which is critical for embodied manipulation. We therefore extend the next-chunk diffusion model in Sec. 2.1 into a multi-view video generator that is conditioned on camera geometry and learns cross-view consistency end-to-end. Given a task prompt $c$, $m$-view observations $O_{1:K}\in\mathbb{R}^{K\times m\times H\times W\times 3}$ with per-view camera intrinsics and extrinsics, we encode each frame into VAE latents and directly predict future multi-view latents $\mathbf{z}_{1:M}\in\mathbb{R}^{M\times m\times H\times W\times C}$, as illustrated in Fig. 2. To make the representation view-aware, we compute a per-view ray-direction map that encodes the camera intrinsics and extrinsics and concatenate it channel-wise with the image latents before the diffusion backbone, following ray-based conditioning chen2023ray ; sajjadi2022scene . Cross-view geometric coherence is modeled using attention along the view dimension: we reshape features to attend across views at corresponding spatial locations and apply spatial attention that preserves pixel-to-ray alignment. Temporal attention is applied along the time dimension to capture scene dynamics. During training, simulator data provide ground-truth camera parameters; for real-world videos we use estimated extrinsics relative to the robot base. At inference, we use the available extrinsics; when depth is available, we optionally depth-warp observed frames to synthesize auxiliary rendered frames that better match the multi-view training conditions.
@@
Figure 3: The pipeline for EnerVerse as a data engine. Observation images from multiple cameras and rendered images are processed by the multi-view video generator to produce denoised videos. These videos, along with their camera poses, are used in 4DGS for 4D scene reconstruction. The reconstructed 3D content is rendered to generate high-precision images. These high-quality rendered images are iteratively refined and fed back into the pipeline.
Real-World Data Flywheels. To reduce reliance on costly fully calibrated multi-camera capture and to narrow sim-to-real gaps, we introduce an offline data flywheel, e.g., EnerVerse-D, that leverages sparse real observations to bootstrap geometry-consistent multi-view videos and progressively reduce the gap, as shown in Fig. 3. After pretraining the 4D base model (EnerVerse-G) as above, we fine-tune it to accept sparse multi-view inputs in the 4D latent space $\mathbb{R}^{C\times V\times T\times H\times W}$. Specifically, among $m$ views, at least $n\ll m$ robot-mounted cameras provide complete observation sequences; for these observed views we skip noise injection and use their clean latents as conditioning, while we apply the standard noisy-to-denoised diffusion to unobserved target views. Given sparse observations (e.g., one full video), their camera parameters, and a task prompt, the model produces denoised predictions for the missing views. We then reconstruct a 4D scene using 4D Gaussian Splatting (4DGS) from the union of observed and generated multi-view videos and their poses. The 4DGS representation is rendered to all target views to obtain higher-fidelity, geometry-consistent frames. These renders can be re-noised and fed back through the multi-view generator, followed by another 4DGS optimization step, yielding an iterative loop that progressively reduces noise, improves reconstruction accuracy, and tightens cross-view consistency. As the loop accumulates real multi-view episodes, we further fine-tune the multi-view generator on the collected data.